# Beyond Nonconvexity: A Universal Trust-Region Method with New Analyses

### v3

@@ Abstract
The trust-region (TR) method is renowned historically for its robustness in nonconvex problems and extraordinary numerical performance, but the study of its performance in convex optimization is somehow limited. This paper complements the existing literature by presenting a universal trust-region method that simultaneously incorporates the quadratic regularization and ball constraint. In particular, we introduce a novel descent property tailored for trust-region-type algorithms, enabling us to unify and streamline the analysis for both convex and nonconvex optimization. Our method exhibits an iteration complexity of $\tilde{O}(\epsilon^{-3/2})$ to find an $\epsilon$-approximate second-order stationary point for nonconvex optimization. Meanwhile, the analysis reveals that the universal method attains an $O(\epsilon^{-1/2})$ complexity bound for convex optimization. Finally, we develop an adaptive universal method to address practical implementations. The numerical results show the effectiveness of our method in both nonconvex and convex problems.
@@ 1 Introduction
$\min_{d\in\mathbb{R}^{n}}$ $m_{k}(d):=\nabla f(x_{k})^{T}d+\frac{1}{2}d^{T}\nabla^{2}f(x_{k})d,$ (1.1)
@@ 2 The Universal Trust-Region Method
$\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq M\|x-y\|\quad\forall x,y\in\mathbb{R}^{n}.$ (2.5)
@@ 7               Go to Line 3;
For some $10<\xi<$ 1, $\kappa>0$, the step $d_{k}$ either decreases the value of the objective function or decreases the gradient norm sufficiently, that is for each iteration $k$,
@@ 2.2 Basic Properties of the Method
$\qquad-\frac{1}{2}d_{k}^{T}\left(\nabla^{2}f(x_{k})+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I+\lambda_{k}I\right)d_{k}+\frac{M}{6}\|d_{k}\|^{3}$
@@
$\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)\cdot r_{k}^{3}\geq\frac{1}{81\sqrt{M}}\quad\text{and}\quad\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}=\frac{1}{6}.$
@@ 3.2 Minimizing Convex Functions
where $10<\xi<$ 1 is the same as that defined in Property 2.2.
@@ 3.3 Local Convergence
$\nabla f(x^{*})=0,\quad\nabla^{2}f(x^{*})\succeq\mu I\succ 0.$ (3.11)
@@ 15       Update $x_{k+1}=x_{k}+d_{k}^{(j)}$, $\rho_{k+1}=\max\{\rho_{\min},\rho_{k}^{(j)}/\gamma_{2}\}$;
{$f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{3/2}\text{ or }\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\epsilon^{3/2},\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon,\end{aligned}\right.$ (4.1)
@@
$&\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ &\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\epsilon+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon.\end{aligned}\right.$ (4.2)
@@
$-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$ $\sigma_{k}^{(j)}=\rho_{k}^{(j)},\ r_{k}^{(j)}=1/4\rho_{k}^{(j)}$
@@ 4.2 Global Convergence
$\mathcal{F}_{p_{s}}$ =$=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\max\left\{\|\nabla f(x_{k})\|^{3/2},\epsilon^{3/2}\right\}\right\},$ (4.5)
$\mathcal{G}_{p_{s}}$ =$=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\},$
@@
$\mathcal{F}_{p_{s}}$ =$=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\|\nabla f(x_{k})\|^{3/2}\right\},$
$\mathcal{G}_{p_{s}}$ =$=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\}.$
@@ 5 Numerical Experiments
In this section, we present numerical experiments. We implement the adaptive UTR (Algorithm 2) in Julia programming language. All experiments are performed on a MacBook Pro with Apple M2 Chip and 24GB LPDDR5 Memory. To enable efficient routines for trust-region subproblems, we implement two options. The first option utilizes the standard Cholesky factorization [2, Algorithm 4.3] and uses a hybrid bisection and Newton method to find the dual variable [32, 33]. When using this option, we name the method after UTR. The second option is an indirect method (so it is referred to as iUTR) by Krylov subspace iterations, which is consistent with the open-source implementation [34] of classical trust-region method and adaptive cubic regularized Newton method. Motivated from Curtis and Wang [25] and Cartis et al. [1, Chapter 10], we use the Lanczos method with inexactness of subproblem solutions.
@@
where $a_{i}\in\mathbb{R}^{n},~{}b_{i}\in\{-1,1\}$, $i=1,2,\cdots,N.$ We set $\gamma=10^{-8}$ so that the Newton steps may fail at degenerate Hessians. Since the problem is convex, we focus on comparisons with the adaptive Newton method with cubics (ArC, Cartis et al. [16]) and variants of the regularized Newton method [22]. We implement the regularized Newton method (RegNewton) with fixed regularization $\sigma_{k}\in\{10^{-3},5\times 10^{-4}\}$ and an adaptive version following Mishchenko [22, Algorithm 2.3] that is named after RegNewton-AdaN+.
@@
A final experiment was conducted for matrix completion problems as a comparison to a recent TR method, CAT [20]. Briefly, the matrix completion task involves reconstructing a partially observed matrix $D\in\mathbb{R}^{n_{1}\times n_{2}}$. The goal is to recover the original dataset $D$ by a partially observed set $\Omega=\{(i,j)\}$ and the low-rank representation $D=PQ^{T}$, where $P\in\mathbb{R}^{n_{1}\times r}$ and $Q\in\mathbb{R}^{n_{2}\times r}$ with $r\ll n_{1}$ and $r\ll n_{2}$. When predicting the power consumption, $n_{1},n_{2}$ are number of features and data points, respectively. Allow a baseline regression model:
@@
where $\mu$ represents the average of all observed measurements, $r_{i},c_{j}$ captures the deviations from $i$ and $j$. We focus on the following regularized problem:
@@
to fit the observed data under (5.2). The dataset used for this study is sourced from Ausgrid [20] that deals with the matrix size of $48\times 30$, $r=9$, all algorithms are terminated at $\|\nabla f(x_{k})\|\leq 10^{-7}$. We vary $\lambda\in[10^{-2},10^{-3},10^{-4}]$ to test the robustness under small regularizations. The results are presented in Table 4. While smaller $\lambda$ induces harder instances, iUTR has the best performance among competing algorithms in all instances. Unfortunately, the CAT method only converges to inaccuracy solution with $\|\nabla f(x_{k})\|\leq 10^{-5}$; hence we only include its results when $\lambda=10^{-2}$.
@@ Appendix A Technical Proofs
    $-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}.$

### v4

# Yuntian Jiang School of Information Management and Engineering

@@ Abstract
‘The trust-region (TR) method is renowned historically for its robustness in nonconvex problems and extraordinary numerical performance, but the study of its performance in convex optimization is somehow limited. This paper complements the existing literature by presenting a universal trust-region method that simultaneously incorporates the quadratic regularization and ball constraint. In particular, we introduce a novel descent property tailored for trust-region-type algorithms, enabling us to unify and streamline the analysis for both convex and nonconvex optimization. Our method exhibits an iteration complexity of $\tilde{O}(\epsilon^{-3/2})$ to find an $\epsilon$-approximate second-order stationary point for nonconvex optimization. Meanwhile, the analysis reveals that the universal method attains an $O(\epsilon^{-1/2})$ complexity bound for convex optimization. Finally, we develop an adaptive universal method to address practical implementations. The numerical results show the effectiveness of our method in both nonconvex and convex problems.
@@ 1 Introduction
$\min_{d\in\mathbb{R}^{n}}$ $m_{k}(d)=\nabla f(x_{k})^{T}d+\frac{1}{2}d^{T}\nabla^{2}f(x_{k})d,$ (1.1)
@@ 2 The Universal Trust-Region Method
$\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq M\|x-y\|\qquad\forall x,y\in\mathbb{R}^{n}.$ (2.5)
@@ 7           Go to Line 3;
For some $0<\xi<1$, $\kappa>0$, the step $d_{k}$ either decreases the value of the objective function or decreases the gradient norm sufficiently, that is for each iteration $k$,
@@ 2.2 Basic Properties of the Method
$\hskip 17.00024pt-\frac{1}{2}d_{k}^{T}\left(\nabla^{2}f(x_{k})+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I+\lambda_{k}I\right)d_{k}+\frac{M}{6}\|d_{k}\|^{3}$
@@
$\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)\cdot r_{k}^{3}\geq\frac{1}{81\sqrt{M}}\qquad\text{and}\qquad\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}=\frac{1}{6}.$
@@ 3.2 Minimizing Convex Functions
where $0<\xi<1$ is the same as that defined in Property 2.2.
@@ 3.3 Local Convergence
$\nabla f(x^{*})=0,\qquad\nabla^{2}f(x^{*})\succeq\mu I\succ 0.$ (3.11)
@@ 15     Update $x_{k+1}=x_{k}+d_{k}^{(j)}$, $\rho_{k+1}=\max\{\rho_{\min},\rho_{k}^{(j)}/\gamma_{2}\}$;
$\small\left\{\begin{aligned} f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{3/2}\text{ or }\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\epsilon^{3/2},\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon,\end{aligned}\right$. (4.1)
@@
$\small\left\{\begin{aligned} &\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ &\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\epsilon+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon.\end{aligned}\right$. (4.2)
@@
$\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$ $\sigma_{k}^{(j)}=\rho_{k}^{(j)},\ r_{k}^{(j)}=1/4\rho_{k}^{(j)}$
@@ 4.2 Global Convergence
$\mathcal{F}_{p_{s}}$ $=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\max\left\{\|\nabla f(x_{k})\|^{3/2},\epsilon^{3/2}\right\}\right\},$ (4.5)
$\mathcal{G}_{p_{s}}$ $=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\},$
@@
$\mathcal{F}_{p_{s}}$ $=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\|\nabla f(x_{k})\|^{3/2}\right\},$
$\mathcal{G}_{p_{s}}$ $=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\}.$
@@ 5 Numerical Experiments
In this section, we present numerical experiments. We implement the adaptive UTR (Algorithm 2) in Julia programming language. All experiments are performed on a MacBook Pro with Apple M2 Chip and 24GB LPDDR5 Memory. To enable efficient routines for trust-region subproblems, we implement two options. The first option utilizes the standard Cholesky factorization [28, Algorithm 4.3] and uses a hybrid bisection and Newton method to find the dual variable [34, 35]. When using this option, we name the method after UTR. The second option is an indirect method (so it is referred to as iUTR) by Krylov subspace iterations, which is consistent with the open-source implementation [14] of classical trust-region method and adaptive cubic regularized Newton method. Motivated from Curtis and Wang [11] and Cartis et al. [6, Chapter 10], we use the Lanczos method with inexactness of subproblem solutions.
@@
where $a_{i}\in\mathbb{R}^{n},~b_{i}\in\{-1,1\}$, $i=1,2,\cdots,N.$ We set $\gamma=10^{-8}$ so that the Newton steps may fail at degenerate Hessians. Since the problem is convex, we focus on comparisons with the adaptive Newton method with cubics (ArC, Cartis et al. [4]) and variants of the regularized Newton method [24]. We implement the regularized Newton method (RegNewton) with fixed regularization $\sigma_{k}\in\{10^{-3},5\times 10^{-4}\}$ and an adaptive version following Mishchenko [24, Algorithm 2.3] that is named after RegNewton-AdaN+.
@@
A final experiment is conducted on matrix completion problems motivated by a recent TR method, CAT [19]. The goal is to recover a power consumption matrix $D\in\mathbb{R}^{n_{1}\times n_{2}}$ using a partially observed set $\Omega=\{(i,j)\}$ and the low-rank representation $D=PQ^{T}$, where $P\in\mathbb{R}^{n_{1}\times r}$ and $Q\in\mathbb{R}^{n_{2}\times r}$ with $r\ll n_{1}$ and $r\ll n_{2}$. In the above $n_{1}$ represents the number of measurements taken per day within a 15 mins interval and $n_{2}$ represents the number of days. We allow the same baseline estimate as [19]:
@@
where $\mu$ represents the average of all observed measurements, $r_{i},c_{j}$ captures the deviations during time $i$ and day $j$. We focus on the following regularized problem:
@@
to fit the observed data under (5.2), where $D_{i,j}$ denotes the observed data recorded at time $i$ and day $j$. The dataset used for this study is sourced from Ausgrid [19] that deals with the matrix size of $48\times 30$, $r=9$, all algorithms are terminated at $\|\nabla f(x_{k})\|\leq 10^{-7}$. We vary $\lambda\in[10^{-2},10^{-3},10^{-4}]$ to test the robustness under small regularizations. The results are presented in Table 4. While smaller $\lambda$ induces harder instances, iUTR has the best performance among competing algorithms in all instances. Since the CAT method only converges to $\epsilon$-FOSPs according to its theory, we only include its results when $\lambda=10^{-2}$.
@@ Appendix A Technical Proofs
    $\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}.$